### v3

# Simpler Hamiltonian structure of Gravity in degenerate limit

Abstract
We analyze Hamiltonian gravity in the limit of a vanishing (spatial) triad determinant. In the context of Barbero-Immirzi formulation, the Lorentzian constraints in this limit are found to exhibit a formal similarity to the Euclidean constraints of Ashtekar’s self-dual formulation, modulo the difference that the spatial diffeomorphisms become trivial. Further, we show that the Carrollian limit of gravity emerges as a special case, thus providing it a geometric interpretation independent of either the speed of light or coupling constants.
I Introduction
The obstacles to a canonical quantization of Lorentzian gravity are fairly well-known, irrespective of whether the canonical variables are defined within a metric or a connection formulation. The structure of the Hamiltonian constraint itself, containing both extrinsic and spatial curvature, poses a serious challenge given the inherent operator ordering ambiguity, non-polynomiality and regularization issues. The hunt for a simpler Hamiltonian structure have also led to efforts towards constructing different limits of gravity. It could be worthwhile to mention a few of these here, such as the strong coupling limit [1, 2], $G\rightarrow 0$ limit [3, 4], $G\rightarrow\infty$ limit [5], the topological limit [6] and so on.
In the Hamiltonian theory of Hilbert-Palatini gravity based on the connection formalism, the densitized triad fields emerge as a natural canonical variable (momenta). In a quantum theory, a discrete quantum geometry could be anticipated, where trivial densitized triads could potentially define the lowest eigenstates provided the other (regularized) constraints are also satisfied. Hence, it is worthwhile to explore classical gravity first in a limit where the spatial geometry becomes trivial in the sense of densitized triads and the determinant of the four-metric. Such a limit for gravity has not been explicitly constructed so far.
To this end, in this brief note we explore a limit where the determinant of the spatial triad (and also of the four-metric) vanishes. We implement the limit by defining a local scaling transformation on the canonical variables within the connection formalism.
First, we apply this limit to the Barbero-Immirzi Hamiltonian formulation [7, 8]. We elucidate two different reductions of the constraints based on two different degenerate limits introduced here. In both cases, the resulting Hamiltonian theories reflect remarkable simplications. In the first case, we find that the Lorentzian constraints exhibit a formal similarity to the Euclidean constraints of Ashtekar’s self-dual gravity [9, 10, *ashtekar1], modulo differences regarding the action of spatial diffeomorphism. As an example, we find a solution to this limiting gravity theory in the context of inhomogeneous and anisotropic cosmology. In case of the second limit, we find a Hamiltonian structure whose quantum counterpart would be free of any operator ordering ambiguity. Next, we find that a special case of the degenerate limit corresponds to Carrollian gravity, the latter being originally defined for a vanishingly small speed of light [12, 13, 1, 2, 14, 15].
Apart from exploring possible new limits of Hamiltonian gravity, our analysis could provide a fresh perspective towards a perturbative expansion of gravity around vanishing densitized triad, different from the standard approach of expanding around a fixed background geometry with $G$ being the perturbative parameter.
In the next section, we first set up the relevant notation through a very brief outline of Barbero-Immirzi canonical gravity. The limit to a degenerate tetrad through a set of scaling transformations is then introduced. The implications of this limit for the Hamiltonian theory is discussed, along with an explicit example. In Sec-III we explore a different limit associated with a degenerate tetrad. This is followed by an analysis of the special case of global scaling transformations, the resulting theories being shown to be equivalent to the Carrolian gravity theories. We conclude our analysis with a few relevant remarks. In the Appendix, as a relevant digression, we present a corollary on Birkhoff’s theorem in ‘electric’ Carroll gravity based on the associated Hamiltonian structure.
II Degenerate limit of Barbero-Immirzi gravity
A modern perspective on Hamiltonian gravity based on the connection formulation is provided by the Barbero-Immirzi $SU(2)$ formulation. This is based on a Lagrangian density containing the Hilbert-Palatini and the topological Nieh-Yan density with a constant coefficient [16, 17], namely the Barbero-Immirzi parameter :
${\cal L}(e,\omega)=\frac{1}{2\kappa}\Big{[}ee^{\mu}_{I}e^{\nu}_{J}R_{\mu\nu}^{~{}~{}~{}IJ}(\omega)~{}+~{}\eta\epsilon^{\mu\nu\alpha\beta}\Big{(}D_{\mu}e_{\nu}^{I}D_{\alpha}e_{\beta I}-e_{\mu}^{I}e_{\nu}^{J}R_{\alpha\beta}^{~{}~{}~{}KL}(\omega)\Big{)}\Big{]}$ (1)
Assuming the standard parametrization of the tetrad as presented in Ref.[17, 18], the totally constrained Hamiltonian density reads:
${\cal H}=NH+N^{a}H_{a}+\frac{1}{2}\omega_{t}^{~{}IJ}G_{IJ}.$ (2)
The scalar Lapse $N$, the shift $N^{a}$ and the temporal connection fields $\omega_{t}^{~{}IJ}$ emerge as Lagrange multipliers above, whereas $H\approx 0,~{}H_{a}\approx 0,~{}G_{IJ}\equiv(G_{0i},~{}G_{ij})\approx 0$ denote the Hamiltonian, spatial diffeomorphism, boost and rotation constraints, respectively. Their explicit expressions are presented later. The spatial metric is defined as: $q_{ab}~{}:=~{}e_{a}^{I}e_{bI}$, and the tetrad determinant as: $e:=det(e^{I}_{\mu})=N\sqrt{q}$.
In time gauge ($e_{a}^{0}\approx 0$), which fixes the boost freedom, the number of canonical pairs get reduced and the expressions of the constraints simplify. The resulting symplectric form reads  [^1] The original canonical pair is defined as: $\omega_{a}^{(\eta)IJ}:=\omega_{a}^{IJ}+\frac{\eta}{2}\epsilon^{IJKL}\omega_{aKL},~{}\pi^{a}_{~{}IJ}:=\frac{1}{2\kappa}\epsilon^{abc}\epsilon_{IJKL}e_{b}^{K}e_{c}^{L}$:
$\Omega=\frac{1}{2}\pi^{a}_{~{}IJ}\partial_{t}\omega_{a}^{(\eta)IJ}=E^{a}_{i}\partial_{t}A_{a}^{i},$ (3)
where the conjugate pair is given by: $A_{a}^{i}:=\omega_{a}^{0i}+\frac{1}{2}\epsilon^{ijk}\omega_{a}^{jk},~{}E^{a}_{i}:=\pi^{a}_{~{}0i}=ee^{a}_{i},~{}q=\det E^{a}_{i}:=E$. The connection fields $A_{a}^{i}$ emerge as the $SU(2)$ gauge field. In terms of this canonical pair, the set of constraints read [19, 16]:
$G^{rot}_{i}~{}$ $=$ $~{}-\eta D_{a}(A)E^{a}_{i},$
$H_{a}~{}$ $=$ $~{}E^{b}_{i}F_{ab}^{i}(A),$
$H~{}$ $=$ $~{}\frac{\kappa}{2\sqrt{E}}E^{a}_{i}E^{b}_{j}\Big{[}\eta\epsilon^{ijk}F_{ab}^{k}(A)-(1+\eta^{2})\Big{(}\partial_{[a}\omega_{b]}^{~{}ij}(E)+\omega_{[a}^{~{}ik}(E)\omega_{b]}^{~{}kj}(E)\Big{)}\Big{]},$ (4)
where $D_{a}(A)V^{i}:=\partial_{a}V^{i}-\frac{1}{\eta}\epsilon^{ijk}A_{a}^{j}V^{k}$ and $F_{ab}^{~{}i}(A):=\partial_{[a}A_{b]}^{i}-\frac{1}{\eta}\epsilon^{ijk}A_{a}^{j}A_{b}^{k}$ denote the $SU(2)$ covariant derivative and field-strength, respectively, and the spatial connection components $\omega_{a}^{~{}ij}(E)$ are defined as:
$\omega_{a}^{~{}ij}(E)~{}=~{}\frac{1}{2}E_{a}^{[i}\partial_{b}E^{bj]}~{}+~{}\frac{1}{2}\epsilon^{ijk}E_{a}^{l}\Big{[}\epsilon^{mn(k}E^{l)}_{b}-\epsilon^{mnp}E_{b}^{p}\delta^{kl}\Big{]}\partial_{c}\big{(}E^{c}_{m}E^{b}_{n}\big{)}$ (5)
Limit to degenerate tetrad
In order to implement the limit to vanishing triad determinant, we introduce a local scaling transformation on the triad fields as:
$e_{a}^{i}=\delta_{(a)}e^{{}^{\prime}i}_{a}.$ (6)
The brackets in the spatial label ‘$(a)$’, which can run over the three space coordinates, imply that it is not summed over when repeated. The primed variables are assumed to remain finite in the limit. To emphasize, each of the three vectors for a fixed $(a)$ need not vanish in this limit. The triad determinant $\sqrt{q}=\delta\sqrt{q^{\prime}}$ goes to zero as $\delta=\prod_{a}\delta_{(a)}\rightarrow 0$, which defines the ‘degenerate’ limit. This in turn implies that: $\delta_{(a)}\delta_{(b)}=\frac{\delta}{\delta_{(c)}}\rightarrow 0$ for any $a\neq b$ in the limit. An explicit realization of these properties is presented through an example later.
Under the condition that the symplectic form be preserved, eq.(6) leads to the following transformation of the canonical pairs:
$A_{a}^{i}$ $=$ $\frac{\delta_{(a)}}{\delta}A_{a}^{{}^{\prime}i},~{}E^{a}_{i}=\frac{\delta}{\delta_{(a)}}E^{{}^{\prime}a}_{i}.$ (7)
Note that $E^{a}_{i}\rightarrow 0$ as $\delta\rightarrow 0$. The natural transformation of the spatial derivative which preserves the tetrad one-form is given by: $\partial^{\prime}_{a}=\frac{\partial_{a}}{\delta_{(a)}}$. Throughout this article we assume that the coupling constants do not scale.
Under (7), the transformed constraints are given by:
$G^{rot}_{i}$ $=$ $\epsilon^{ijk}A_{a}^{{}^{\prime}k}E^{{}^{\prime}a}_{j}-\eta\frac{\delta}{\delta_{(a)}}\partial_{a}E^{{}^{\prime}a}_{i}-\eta\partial_{a}\frac{\delta}{\delta_{(a)}};$
$H_{a}$ $=$ $\frac{\delta_{(a)}}{\delta}\Big{[}-\frac{1}{\eta}\epsilon^{ijk}A_{a}^{{}^{\prime}j}A_{b}^{{}^{\prime}k}E^{{}^{\prime}b}_{i}-\frac{\delta}{\delta_{(a)}}E^{{}^{\prime}b}_{i}\partial_{a}A_{b}^{{}^{\prime}i}-\frac{\delta}{\delta_{(b)}}E^{{}^{\prime}b}_{i}\partial_{b}A_{a}^{{}^{\prime}i}$
-- $E^{{}^{\prime}b}_{i}A_{b}^{{}^{\prime}i}\frac{\delta}{\delta_{(a)}}\frac{\delta_{(b)}}{\delta}\partial_{a}\frac{\delta}{\delta_{(b)}}+E^{{}^{\prime}b}_{i}A_{a}^{{}^{\prime}i}\frac{\delta}{\delta_{(b)}}\frac{\delta_{(a)}}{\delta}\partial_{a}\frac{\delta}{\delta_{(a)}}\Big{]},$
$H$ $=$ $\frac{\kappa}{2\delta\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}A_{a}^{{}^{\prime}i}A_{b}^{{}^{\prime}j}~{}+~{}\delta^{2}\Big{(}\frac{1}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}+\omega_{a}^{{}^{\prime}ik}\omega_{b}^{{}^{\prime}kj}\Big{)}+\omega^{{}^{\prime}ij}_{b}\frac{\delta^{2}}{\delta_{(a)}\delta_{(b)}}\partial_{a}\frac{\delta_{(b)}}{\delta}\Big{]}$
In writing the last line above we have used the expression (5) for the spatial connection components. With the assumptions: $\partial_{a}\frac{\delta}{\delta_{(b)}}\rightarrow 0\Rightarrow~{}\frac{\delta}{\delta_{(a)}}\frac{\delta}{\delta_{(b)}}\partial_{c}\frac{\delta_{(b)}}{\delta}\rightarrow 0$ for any $a,b,c$, the spatial derivative terms in the constraints above drop out. Redefining the shift and lapse as: $N^{a}=\frac{\delta}{\delta_{(a)}}N^{{}^{\prime}a},~{}N=\delta N$, we find, as $\delta\rightarrow 0$:
$G^{{}^{\prime}rot}_{i}$ $=$ $\epsilon^{ijk}A_{a}^{{}^{\prime}k}E^{{}^{\prime}a}_{j};$
$H^{\prime}_{a}$ $=$ $\frac{1}{\eta}A_{a}^{{}^{\prime}}G^{{}^{\prime}rot}_{i};$
$H^{\prime}$ $=$ $-\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}A_{a}^{{}^{\prime}i}A_{b}^{{}^{\prime}j}$ (8)
Noting that the $SU(2)$ covariant derivative and the field-strength in the limit are given by: $D^{\prime}_{a}(A^{\prime})E^{{}^{\prime}a}_{i}=-\frac{1}{\eta}\epsilon^{ijk}E^{{}^{\prime}b}_{j}A_{a}^{{}^{\prime}j}E^{{}^{\prime}ak},~{}F_{ab}^{{}^{\prime}~{}i}=-\frac{1}{\eta}\epsilon^{ijk}A^{{}^{\prime}j}_{a}A_{b}^{{}^{\prime}k}$, the above set may be rewritten as:
$G^{{}^{\prime}rot}_{i}$ $=$ $\eta D^{\prime}_{a}(A^{\prime})E^{{}^{\prime}a}_{j};$
$H^{\prime}_{a}$ $=$ $E^{{}^{\prime}b}_{i}F_{ab}^{{}^{\prime}~{}i}(A^{\prime});$
$H^{\prime}$ $=$ $\frac{\kappa\eta}{2\sqrt{E^{\prime}}}\epsilon^{ijk}E^{{}^{\prime}a}_{i}E^{{}^{\prime}b}_{j}F_{ab}^{{}^{\prime}~{}k}(A^{\prime})$ (9)
This form precisely resembles the set of Euclidean constraints of Ashtekar’s self-dual gravity (which could be obtained by setting $\eta=i$ in Barbero-Immirzi formulation). The important difference, however, is that the Lorentzian set here corresponds to a simpler canonical structure, given that the spatial diffeomorphism constraints are trivially satisfied.
Example: Ultralocal physics
The properties of the scaling fields $\delta_{(a)}$ allow us to describe a dynamical spacetime where all the spatial directions need not contract in general as the triad (tetrad) degenerates. As an example of an inhomogeneous and anisotropic universe, let us consider the following spacetime:
$ds^{2}=-dt^{2}+t^{2\alpha(x,y,z)}dx^{2}+t^{2\beta(x,y,z)}dy^{2}+t^{2\gamma(x,y,z)}dz^{2}$
where we assume that: $\alpha+\beta>0,~{}\beta+\gamma>0,~{}\alpha+\gamma>0$. A similar metric was first considered by BKL [20, *bkl1], which could be a solution of Einstein gravity in presence of matter. Assuming that the effect of matter could be neglected in the degenerate limit (which here represents a curvature singularity at $t=0$), we proceed to find the associated solution.
Let us first demonstrate that the assumed generic properties of the scaling transformation are all satisfied. The metric degenerates as $t\rightarrow 0$. This limit could be parametrized as $\epsilon\rightarrow 0$ where $t=\epsilon t^{\prime}$ ($t^{\prime}$ being finite). Inserting this in the metric solution just obtained above, we find:
$\delta_{(x)}=\epsilon^{\alpha(x,y,z)},~{}\delta_{(y)}=\epsilon^{\beta(x,y,z)},~{}\delta_{(z)}=\epsilon^{\gamma(x,y,z)}.$
Thus, the degenerate limit is equivalent to $\delta=\epsilon^{\alpha+\beta+\delta}\rightarrow 0$. Let us note that since $\epsilon^{\sigma}\ln\epsilon\rightarrow 0$ as $\epsilon\rightarrow 0$ for any $\sigma>0$ (as a consequence of L’Hospital’s rule), we have: $\partial_{a}\frac{\delta}{\delta_{(b)}}\rightarrow 0$ for any $a,b$. This in turn implies that $\frac{\delta}{\delta_{(a)}}\frac{\delta_{(b)}}{\delta}\partial_{c}\frac{\delta}{\delta_{(b)}}\rightarrow 0$ for any $a,b,c$. Thus, the scaling transformation satisfies the assumptions made earlier.
The nontrivial components of the momenta are given by (suppressing the primes which are implicit in the variables and constraints):
$E^{x}_{1}=t^{\beta+\gamma},~{}E^{y}_{2}=t^{\gamma+\alpha},~{}E^{z}_{3}=t^{\alpha+\beta}.$
The time evolution of the canonical pair with respect to the Hamiltonian constraint (II) leads to (in the gauge $\omega_{t}^{~{}ij}=0$):
$\dot{E}^{a}_{i}$ $=$ $\left[E^{a}_{i},\int{\cal H}\right]=\frac{N}{\sqrt{E}}E^{[a}_{i}E^{b]}_{j}A_{b}^{j};$
$\dot{A}_{a}^{i}$ $=$ $\left[A_{a}^{i},\int{\cal H}\right]=-\frac{N}{\sqrt{E}}\Big{[}E^{b}_{j}A_{[a}^{i}A_{b]}^{j}+\frac{\sqrt{E}}{2}E_{a}^{i}H\Big{]}$ (10)
From the first equation above, the nontrivial components of the coordinates are obtained as:
$A_{x}^{1}=\alpha t^{\alpha-1},~{}A_{y}^{2}=\beta t^{\beta-1},~{}A_{z}^{3}=\gamma t^{\gamma-1}$
The second evolution equation in (II) for the coordinates, along with the Hamiltonian constraint, imply:
$\alpha\beta+\beta\gamma+\gamma\alpha=0,$
$\alpha(\alpha+\beta+\gamma-1)=0,$
$\beta(\alpha+\beta+\gamma-1)=0,$
$\gamma(\alpha+\beta+\gamma-1)=0.$ (11)
Assuming $\alpha\neq 0,~{}\beta\neq 0,~{}\gamma\neq 0$, these are solved as: $\alpha+\beta+\gamma=1=\alpha^{2}+\beta^{2}+\gamma^{2}$. The rotation constraints are trivially satisfied.
As the above solution reflects, the spatial dependence of these exponents are arbitrary. Thus, the spatial points are uncorrelated and at each point there could be an independent set of Kasner exponents. In other words, the degenerate limit in this context leads to a classical realization of ultralocal physics.
III A different degenerate limit: Hamiltonian structure without ordering ambiguity
In this section, we explore a different scaling transformation to set up the degenerate limit. This is to elucidate the fact that it is possible to define the limit so that the spatial derivative terms do not necessarily disappear from the constraints.
Let us consider the following scaling transformations:
$A_{a}^{i}$ $=$ $\delta_{(a)}A_{a}^{{}^{\prime}i},~{}E^{a}_{i}=\frac{\delta}{\delta_{(a)}}E^{{}^{\prime}a}_{i},$ (12)
The triad transforms as earlier, whereas the gauge field now transform differently. Note that this does not preserve the symplectic form. Let us also demand that the $E^{a}_{i}\partial_{a}$ be preserved (instead of the tetrad one-form), implying that: $\partial_{a}=\frac{\delta_{(a)}}{\delta}\partial^{\prime}_{a}$.
Using the above transformations, the constraints become:
$G^{rot}_{i}$ $=$ $-\eta\frac{\delta}{\delta_{(a)}}\partial_{a}E^{{}^{\prime}a}_{i}+\delta\epsilon^{ijk}A_{a}^{{}^{\prime}k}E^{{}^{\prime}a}_{j}-\eta E^{{}^{\prime}a}_{i}\partial_{a}\frac{\delta}{\delta_{(a)}},$
$H_{a}$ $=$ $\frac{\delta_{(a)}}{\delta}\Big{[}-\frac{\delta}{\eta}\epsilon^{ijk}A_{a}^{{}^{\prime}j}A_{b}^{{}^{\prime}k}E^{{}^{\prime}b}_{i}-\frac{\delta^{2}}{\delta_{(a)}}E^{{}^{\prime}b}_{i}\partial_{a}A_{b}^{{}^{\prime}i}-\frac{\delta^{2}}{\delta_{(b)}}E^{{}^{\prime}b}_{i}\partial_{b}A_{a}^{{}^{\prime}i}$
$+$ $E^{{}^{\prime}b}_{i}A_{b}^{{}^{\prime}i}\frac{\delta}{\delta_{(a)}}\frac{\delta}{\delta_{(b)}}\partial_{a}\delta_{(b)}-E^{{}^{\prime}b}_{i}A_{a}^{{}^{\prime}i}\frac{\delta}{\delta_{(b)}}\frac{\delta}{\delta_{(a)}}\partial_{b}\delta_{(a)}\Big{]},$
$H$ $=$ $\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}A_{a}^{{}^{\prime}i}A_{b}^{{}^{\prime}j}~{}+~{}(1+\eta^{2})\delta^{2}\Big{(}\frac{1}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}+\omega_{a}^{{}^{\prime}ik}\omega_{b}^{{}^{\prime}kj}\Big{)}~{}+~{}(1+\eta^{2})\omega^{{}^{\prime}ij}_{b}\frac{\delta^{2}}{\delta_{(a)}\delta_{(b)}}\partial_{a}\frac{\delta_{(b)}}{\delta}\Big{]}$
The properties of the scaling transformation imply that the terms involving derivatives of the scaling fields vanish in the limit, leading to:
$G^{{}^{\prime}rot}_{i}$ $=$ $-\eta D^{\prime}_{a}E^{{}^{\prime}a}_{i},$
$H^{\prime}_{a}$ $=$ $0,$
$H^{\prime}$ $=$ $-\frac{\kappa(1+\eta^{2})}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\bar{R}_{ab}^{{}^{\prime}ij}(E^{\prime}),$
where $D^{\prime}_{a}E^{{}^{\prime}a}_{i}=\partial_{a}E^{a}_{i}$ and $\bar{R}_{ab}^{{}^{\prime}ij}(E^{\prime})=\partial_{[a}\omega_{b]}^{{}^{\prime}ij}(E^{\prime})+\omega_{[a}^{{}^{\prime}ik}(E^{\prime})\omega_{b]}^{{}^{\prime}kj}(E^{\prime})$ is the spatial three-curvature. In obtaining the above, we have redefined the Lagrange multipliers as: $\omega_{t}^{~{}ij}=\frac{1}{\delta}\omega_{t}^{{}^{\prime}ij},~{}N^{a}=\frac{\delta}{\delta_{(a)}}N^{{}^{\prime}a}$.
This reflects another intriguing limit, where the diffeomorphism becomes trivial as in the earlier section. The rotation and Hamiltonian constraints above depend only on the conjugate momenta $E^{a}_{i}$. The quantization of this theory would be free of any operator ordering ambiguity problem.
IV A Connection between degenerate and Carroll limit
In this section, we demonstrate that the degenerate limit of Hilbert-Palatini gravity contains Carroll gravity as a special case. The latter had originally been formulated as a truncated Hamiltonian theory in the limit $c\rightarrow 0$. Recently, both the ‘electric’ and ‘magnetic’ forms of this limit was shown to originate from the Hilbert-Palatini action, and the associated connection-based Hamiltonian formulations were also presented [15].
Using the same parametrization as earlier, the Hilbert-Palatini constraints are given by:
$G^{rot}_{i}~{}$ $=$ $~{}\epsilon^{ijk}K_{a}^{k}E^{a}_{j},$
$H_{a}~{}$ $=$ $~{}E^{b}_{i}\partial_{[a}K_{b]}^{i}-K_{a}^{i}\partial_{b}E^{b}_{i},$
$H~{}$ $=$ $~{}-~{}\frac{\kappa}{2\sqrt{E}}E^{[a}_{i}E^{b]}_{j}\Big{[}K_{a}^{i}K_{b}^{j}~{}+~{}\partial_{a}\omega_{b}^{~{}ij}(E)+\omega_{b}^{~{}ik}(E)\omega_{b}^{~{}kj}(E)\Big{]}$ (13)
where $K_{a}^{i}:=\omega_{a}^{~{}0i}$. The analogue of the scaling transformation (7) on conjugate pairs is given by:
$K_{a}^{i}$ $=$ $\frac{\delta_{(a)}}{\delta}K_{a}^{{}^{\prime}i},~{}E^{a}_{i}=\frac{\delta}{\delta_{(a)}}E^{{}^{\prime}a}_{i}$ (14)
Note that the transformation of the derivatives and of $\omega_{a}^{{}^{\prime}ij}(E)$ remain the same as in section-II. Let us now find the reduced set of constraints in the degenerate limit, where the scaling transformation is global: $\delta_{(a)}\equiv const.$
IV.1 Electric Carroll gravity
Under (14), the rotation constraint under the scaling remain form-invariant:
$G^{rot}_{i}=\epsilon^{ijk}K_{a}^{{}^{\prime}k}E^{{}^{\prime}a}_{j}:=G^{{}^{\prime}rot}_{i}$ (15)
The spatial diffeomorphism constraints transform as:
$H_{a}=\delta_{(a)}\Big{[}\frac{1}{\delta_{(a)}}E^{{}^{\prime}b}_{i}\partial_{a}K_{b}^{{}^{\prime}i}-\frac{1}{\delta_{(b)}}E^{{}^{\prime}b}_{i}\partial_{b}K_{a}^{{}^{\prime}i}-\frac{1}{\delta_{(b)}}K_{a}^{{}^{\prime}i}\partial_{b}E^{{}^{\prime}b}_{i}\Big{]}$
Redefining the shift as: $N^{a}:=\frac{1}{\delta_{(a)}}N^{{}^{\prime}a}$, we find:
$H^{\prime}_{a}=E^{{}^{\prime}b}_{i}\partial^{\prime}_{a}K_{b}^{{}^{\prime}i}-E^{{}^{\prime}b}_{i}\partial^{\prime}_{b}K_{a}^{{}^{\prime}i}-K_{a}^{{}^{\prime}i}\partial^{\prime}_{b}E^{{}^{\prime}b}_{i}$ (16)
Finally, the Hamiltonian constraint becomes:
$H=-\frac{\kappa}{2\delta\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}K_{a}^{{}^{\prime}i}K_{b}^{{}^{\prime}j}+~{}\delta^{2}\Big{(}\frac{1}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}+\omega_{a}^{{}^{\prime}ik}\omega_{b}^{{}^{\prime}kj}\Big{)}\Big{]}$
As $\delta\rightarrow 0$, we obtain:
$H^{\prime}=-\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}K_{a}^{{}^{\prime}i}K_{b}^{{}^{\prime}j}$ (17)
with the redefinition: $N:=\delta N^{\prime}$.
Notice that the Hamiltonian structure above corresponds precisely to the ‘electric’ Carroll limit in connection formulation [15]. The additional factor $\frac{1}{\sqrt{E^{\prime}}}$ in the Hamiltonian constraint $H$ here compared to ref.[15] originates due to the use of a scalar lapse.
IV.2 Magnetic Carroll gravity
Next, we consider a global scaling analogous to eq.(12), which provides another way to implement the degenerate limit:
$K_{a}^{i}$ $=$ $\delta_{(a)}K_{a}^{{}^{\prime}i},~{}E^{a}_{i}=\frac{\delta}{\delta_{(a)}}E^{{}^{\prime}a}_{i}.$ (18)
Following the procedure already elucidated in section-III, we find:
$G^{rot}_{i}=\delta\epsilon^{ijk}K_{a}^{{}^{\prime}k}E^{{}^{\prime}a}_{j},$
$H_{a}=\delta_{(a)}\Big{[}\frac{\delta}{\delta_{(a)}}E^{{}^{\prime}b}_{i}\partial_{a}K_{b}^{{}^{\prime}i}-\frac{\delta}{\delta_{(b)}}E^{{}^{\prime}b}_{i}\partial_{b}K_{a}^{{}^{\prime}i}-\frac{\delta}{\delta_{(b)}}K_{a}^{{}^{\prime}i}\partial_{b}E^{{}^{\prime}b}_{i}\Big{]}$
$H=-\frac{\kappa}{2\delta\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}\frac{\delta}{\delta_{(a)}}\partial_{a}\omega_{b}^{{}^{\prime}ij}(E^{\prime})+\omega_{a}^{{}^{\prime}ik}(E^{\prime})\omega_{b}^{{}^{\prime}kj}(E^{\prime})~{}+~{}\delta^{2}K_{a}^{{}^{\prime}i}K_{b}^{{}^{\prime}j}\Big{]}$
Using the same redefinition of the Lagrange multipliers as in sec-III, we find that the rotation and spatial diffeomorphism constraints remain form-invariant, whereas the Hamiltonian constraint reduces to:
$H^{\prime}$ $=$ $-\frac{\kappa}{2\sqrt{E^{\prime}}}E^{{}^{\prime}[a}_{i}E^{{}^{\prime}b]}_{j}\Big{[}\partial^{\prime}_{a}\omega_{b}^{{}^{\prime}ij}(E^{\prime})+\omega_{a}^{{}^{\prime}ik}(E^{\prime})\omega_{b}^{{}^{\prime}kj}(E^{\prime})\Big{]}$
This set of constraints resembles the Hamiltonian form of the ‘magnetic’ Carroll gravity [22, 14] in the connection formulation [15].
V Conclusions
Here we have constructed a limit of Hamiltonian gravity theory corresponding to vanishing densitized triad. This is equivalent to the limit to a vanishing metric determinant. This limit is relevant in the context of connection-triad formulation of canonical quantum gravity.
Here we unravel two different limits of the Barbero-Immirzi formulation. In both cases, the spatial diffeomorphism constraints are satisfied trivially. Further, the respective Hamiltonian constraints exhibit significant simplifications. In the case of the first limit, an ultralocal behaviour (absence of space derivatives) emerges. A similar behaviour is also encountered in other contexts, for instance, in studies of the BKL limit to singularity. However, let us also emphasize that the associated Hamiltonian structure here is quite different from the one corresponding to a BKL behaviour [23, *sloan1]. A canonical quantization of this degenerate limit may be taken up based on either the triad or connection representation. For the second limit, a triad representation would be more appropriate since the constraints in this case depend only on the densitized triad and its derivatives. It could be interesting to explore the possible (kinematical) consequences for quantum geometry, e.g. the area operator in such limits. A study of the possible perturbative theories around this limit could also be worthwhile.
Our analysis also provides a different interpretation to Carroll gravity in terms of the spatial metric determinant. This may be contrasted with the original interpretation in terms of the $c\rightarrow 0$ limit, or a strong gravitational coupling. The question as to which one is to be preferred is purely optional.
Acknowledgements.
This work is supported (in part) by the MATRICS project grant MTR/2021/000008, SERB, Govt. of India.
Appendix A A Corollary on the violation of Birkhoff’s theorem in electric Carroll gravity
As a relevant extension to our discussion of Carroll gravity in this general context, we shall elucidate an important implication of the limiting Hamiltonian theory in the context of spherical symmetry.
Let us first consider ‘electric’ Carrollian limit of gravity defined by the constraints (15), (16) and (17).
Given the standard ansatz for a spherically symmetric spacetime:
$ds^{2}=-e^{\mu(t,r)}dt^{2}+e^{\lambda(t,r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$
the canonical momenta $E^{a}_{i}$ are given by:
$E^{r}_{1}=r^{2}\sin\theta,~{}E^{\theta}_{2}=e^{\frac{\lambda}{2}}r\sin\theta,~{}E^{\phi}_{3}=e^{\frac{\lambda}{2}}r.$ (20)
From their time evolution with respect to the total Hamiltonian, the only nontrivial canonical coordinate reads:
$K_{r}^{1}:=\omega_{r}^{~{}01}=\frac{1}{2}\dot{\lambda}e^{\frac{\lambda-\mu}{2}}.$ (21)
All the constraints are satisfied trivially, except the radial component of the spatial diffeomorphism, which implies:
$Q_{r}^{1}=0\Rightarrow\lambda=\lambda(r)$ (22)
Hence, $\lambda$ could be any arbitrary radial function, whereas the lapse (or, $\mu(t,r)$) is left undetermined.
This demonstrates that the Birkhoff’s theorem is violated by the ‘electric’ limit of Carrollian gravity.
For the other possible Carrollian case, namely the so-called ‘magnetic’ limit, it is straightforward to show that the limiting theory reproduces the Schwarzschild geometry as the unique solution, thus satisfying the Birkhoff’s theorem.